Solve the given initial-value problem

(3y + 2t - 5)dt + (4y+ 3t - 1)dy = 0, y(-1)=2
1
Expert's answer
2020-07-21T18:49:55-0400

D.E is

"(3y + 2t - 5)dt + (4y+ 3t - 1)dy = 0"

Let,

"y=y'+k\&t=t'+h"

Such that

"2h+3k-5=0\\3h+4k-1=0"

Thus, "(h,k)=(13,-7)" and actual equation reduced to homogeneous equation such that

"\frac{dy'}{dt'}=-\dfrac{2t'+3y'}{3t'+4y'}"

Let,

"y'=vt'\implies v+t'\frac{dv}{dt'}=-\dfrac{2+3v}{3+4v}\\
\implies \frac{dt'}{t'}=-\dfrac{4v+3}{4v^2+6v+2}dv\\
\implies \ln|t'|=-\frac{1}{2}\ln|(2v^2+3v+1)|\\
\implies c t'^2=\frac{1}{2v^2+3v+1}\\
\implies ct'^4=\frac{1}{2y'^2+3y't'+t'^2}"

But,

"y'=y+7\&t'=t-13"

Thus,

"c(t-13)^4=\frac{1}{2(y+7)^2+3(y+7)(t-13)+(t-13)^2}"

As, "y(-1)=2" ,thus

"c=\frac{1}{14^4(-20)}"

Then the solution will be

"-\frac{1}{20\cdot14^4}=\frac{1}{2(y+7)^2+3(y+7)(t-13)+(t-13)^2}"